# Green’s functions, propagation invariants, reciprocity theorems, wave-field representations and propagator matrices in 2D time-dependent materials

### v1

@@ 5.1 Unified wave equation in wave-vector time domain
with time-dependent material parameters $\alpha(t)$ and $\beta(t)$. The second order wave equation (9) transforms to
@@ 6.1 Conservation of net field-momentum density
$\check{M}({\bf k},t)=\Re(\check{U}^{*}\check{\bf V})\cdot{\bf n},$ (47)
@@
Propagation invariants in a time-dependent material are wave-field-related quantities that remain constant over time. We start by showing that for a source-free, time-dependent material, the net field-momentum density defined in equation (47) is a propagation invariant. To this end we analyze $\partial_{t}\check{M}({\bf k},t)=\frac{1}{2}\partial_{t}(\check{U}^{*}\check{\bf V}+\check{\bf V}^{*}\check{U})\cdot{\bf n}$. Applying the product rule for differentiation, using equations (39) and (40) for the source-free situation and the definition ${\bf n}={\bf k}/|{\bf k}|$, we obtain
$\partial_{t}\check{M}({\bf k},t)$ $=$ $\frac{1}{2}\Bigl{(}(\partial_{t}\check{U})^{*}\check{\bf V}+\check{U}^{*}\partial_{t}\check{\bf V}+(\partial_{t}\check{\bf V})^{*}\check{U}+\check{\bf V}^{*}\partial_{t}\check{U}\Bigr{)}\cdot{\bf k}/|{\bf k}|$
$=$ $\frac{1}{2}\biggl{(}\Bigl{(}\frac{1}{\beta}i{\bf k}\cdot\check{\bf V}^{*}\Bigr{)}\check{\bf V}-\check{U}^{*}\Bigl{(}\frac{1}{\alpha}i{\bf k}\check{U}\Bigr{)}+\Bigl{(}\frac{1}{\alpha}i{\bf k}\check{U}^{*}\Bigr{)}\check{U}-\check{\bf V}^{*}\Bigl{(}\frac{1}{\beta}i{\bf k}\cdot\check{\bf V}\Bigr{)}\biggr{)}\cdot{\bf k}/|{\bf k}|=0.$
@@
$\check{\jmath}({\bf k},t)=\frac{c^{2}(t)}{2}\check{M}({\bf k},t),$ (49)
which implies that for a time-dependent medium with piecewise continuous parameters the net power-flux density is not a propagation invariant, similar as noted in section 3.2 for $j({\bf x},t)$ over a single time boundary.
@@ 6.2 General propagation invariants
Next, we derive more general propagation invariants. We consider two mutually independent states (i.e., solutions of equations (39) and (40)), which we distinguish with subscripts $A$ and $B$. We analyze the quantities $\partial_{t}(\check{U}_{A}\check{\bf V}_{B}-\check{\bf V}_{A}\check{U}_{B})\cdot(i{\bf k})$ and $\partial_{t}(\check{U}_{A}^{*}\check{\bf V}_{B}+\check{\bf V}_{A}^{*}\check{U}_{B})\cdot(i{\bf k})$. The latter can be seen as a generalization of $2\partial_{t}\check{M}({\bf k},t)=\partial_{t}(\check{U}^{*}\check{\bf V}+\check{\bf V}^{*}\check{U})\cdot{\bf n}$, with subscripts $A$ and $B$ added and ${\bf n}={\bf k}/|{\bf k}|$ replaced by $i{\bf k}$ (the latter replacement facilitates the inverse Fourier transformations in later sections). Applying the product rule for differentiation, using equations (39) and (40) for states $A$ and $B$, yields
@@ 7.2 Source-receiver reciprocity
This is the counterpart of the well-known source-receiver reciprocity relation ${\cal G}({\bf x}_{B},{\bf x}_{A},t,0)={\cal G}({\bf x}_{A},{\bf x}_{B},t,0)$ for an inhomogeneous time-independent medium, which states that the source and receiver positions can be interchanged (de Hoop, 1995; Morse & Feshbach, 1953). Equation (58) states that for a homogeneous time-dependent medium the source and receiver times can be interchanged if, simultaneously, the causal response is replaced by the acausal response. In the derivation we assumed that $t_{A}$ and $t_{B}$ lie both between $t_{b}$ and $t_{e}$. If we take $t_{b}\to-\infty$ and $t_{e}\to\infty$, then equations (56) – (58) hold for arbitrary $t_{A}$ and $t_{B}$. Note that for $t_{A}<t_{B}$ these expressions reduce to the trivial relation $0=0$.
@@ 7.3 Field-momentum density balance
$\check{M}({\bf k},t)\Bigr{|}_{t_{b}}^{t_{e}}$ $=$ $\Re\int_{t_{b}}^{t_{e}}\bigl{(}\check{a}^{*}\check{\bf V}+\check{\bf b}^{*}\check{U}\bigr{)}\cdot{\bf n}{\rm d}t.$ (60)
@@ 8.2 Wave field representation for the case of a single time boundary
$\check{U}({\bf k},t^{\prime})=-\Bigl{[}\check{\cal G}({\bf k},{\bf 0},t^{\prime},t)\check{V}_{\rm d}({\bf k},t)+\beta_{2}\{\partial_{t}\check{\cal G}_{2}({\bf k},{\bf 0},t^{\prime},t)\}\check{U}({\bf k},t)\Bigr{]}_{t=t_{b}},\quad\mbox{for}\quad t^{\prime}>t_{b}.$ (67)
@@ 9 Green’s function retrieval
$\beta(t_{b})\Bigl{[}\check{\cal G}^{*}({\bf k},{\bf 0},t_{A},t)\partial_{t}\check{\cal G}({\bf k},{\bf 0},t_{B},t)-\{\partial_{t}\check{\cal G}^{*}({\bf k},{\bf 0},t_{A},t)\}\check{\cal G}({\bf k},{\bf 0},t_{B},t)\Bigr{]}_{t=t_{b}}.$ (81)
@@ 10 Matrix-vector wave equation
The subscript ${\rm d}$ in ${\bf q}_{\rm d}$, ${\bf A}_{\rm d}$ and ${\bf d}_{\rm d}$ is used to distinguish these quantities from those in the 1D situation where, for example, wave field vector ${\bf q}$ contains the scalar wave fields $U$ and $V$ (Wapenaar et al., 2024). Equation (92) holds for a homogeneous time-dependent medium with continuous parameters $\alpha(t)$ and $\beta(t)$. For a medium with piecewise continuous parameters, ${\bf q}_{\rm d}({\bf x},t)$ remains continuous at those time instants where $\alpha(t)$ and $\beta(t)$ are discontinuous. Equation (92) is the counterpart of a matrix-vector wave equation for an inhomogeneous time-independent material (Corones, 1975; Kosloff & Baysal, 1983; Fishman & McCoy, 1984; Ursin, 1983; Wapenaar & Berkhout, 1986; Løseth & Ursin, 2007).
@@ 11.1 Propagator matrix in space-time domain
As an illustration, we consider a piecewise constant material consisting of five time-independent slabs, with propagation velocities of 2000, 2000, 1200, 2500 and 1400 m/s, respectively. The parameter $\beta$ is taken constant. The duration of each time slab is 25 $\mu$s. We choose $t_{0}$ between the first and second time slab, i.e., $t_{0}=25\,\mu$s. Figure 4(a) shows a $z,t$-diagram of the propagator element $W_{\rm d}^{U,V}({\bf x},t,t_{0})$ (convolved with a spatial wavelet) for $x=0$ and $t_{0}=25\,\mu$s. The green line at $t_{0}=25\,\mu$s indicates the initial condition of equation (99), which for the considered off-diagonal element implies $W_{\rm d}^{U,V}({\bf x},t_{0},t_{0})=0$. In the first time slab right of this green line, the propagator consists (for $x=0$) of upward and downward propagating direct waves. At each time boundary (indicated by the dashed blue lines) these waves split into upward and downward propagating transmitted and reflected waves. This is a manifestation of the recursive character of the propagator, as formulated by equation (102).
Figure 4: Cross-sections at $x=0$ of (a) the propagator element $W_{\rm d}^{U,V}({\bf x},t,t_{0})$ and (b) the Green’s function ${\cal G}({\bf x},{\bf 0},t,t_{0})$ (both convolved with a spatial wavelet) for a piecewise constant time-dependent material. Movie of (a) available at https://www.keeswapenaar.nl/TimeMaterial/Wuv.mp4.
@@ 11.4 Relations between the Green’s function and the propagator matrix
Comparing this with equation (30) for ${\bf x}_{0}={\bf 0}$, with $c(t)$ defined in equation (10), we conclude that $H(t-t_{0})W_{\rm d}^{U,V}({\bf x},t,t_{0})$ indeed obeys the same wave equation as $-{\cal G}({\bf x},{\bf 0},t,t_{0})$. This completes the proof of equation (119). Figure 4(b) shows a $z,t$-diagram of the Green’s function ${\cal G}({\bf x},{\bf 0},t,t_{0})$ (convolved with a spatial wavelet) for $x=0$ and $t_{0}=25\,\mu$s. It has been obtained from $W_{\rm d}^{U,V}({\bf x},t,t_{0})$ via equation (119). The green line at $t_{0}=25\,\mu$s indicates the causality condition of equation (31).
@@ 12 Conclusions
We used the reciprocity theorem of the space-convolution type to derive a source-receiver reciprocity relation for the Green’s function of a time-dependent material: the causal Green’s function between a source at $t_{B}$ and a receiver at $t_{A}$ is identical to the acausal Green’s function between a sink at $t_{A}$ and a receiver at $t_{B}$. We also used the reciprocity theorem of the space-convolution type to derive a general wave-field representation for a homogeneous time-dependent material. Similar as its counterpart for an inhomogeneous time-independent material, it forms the basis for the analysis of wave scattering problems. As an example, we used this representation to quantitatively explain the behaviour of the Green’s function after a single time boundary. We used the reciprocity theorem of the space-correlation type to derive a representation for Green’s function retrieval in a homogeneous, piecewise continuous time-dependent material.
@@ Acknowledgements
The author thanks Johannes Aichele and Dirk-Jan van Manen for the inspiring discussions about waves in time-dependent materials.
@@ Appendix A Inverse Fourier Transform of wave-vector time domain Green’s function
The integral over $\theta$ is independent of $\phi$, so we choose $\phi=0$. Furthermore, this integral from 00 to $2\pi$ is twice the real part of the integral from 00 to $\pi$. Hence
@@ Appendix B Propagator matrix for a time-independent slab
For the special case of a time-independent slab between $t_{n-1}$ and $t_{n}$, with parameters $\alpha_{n}$ and $\beta_{n}$, the solution of equation (103), with boundary condition (104) (with $t_{0}$ replaced by $t_{n-1}$ and $t$ by $t_{n}$), reads
@@
with $\check{\cal G}({\bf k},{\bf 0},t_{n},t_{n-1})$ given in equation (45) for ${\bf x}_{0}={\bf 0}$, $t_{0}=t_{n-1}$ and $t=t_{n}$. Similarly, we obtain an explicit expression for $W_{\rm d}^{U,V}({\bf x},t_{n},t_{n-1})$ in the space-time domain by substituting equation (129) for ${\bf x}_{0}={\bf 0}$ into equation (124). This gives

### v2

@@ 5.1 Unified wave equation in wave-vector time domain
with time-dependent material parameters $\alpha(t)$ and $\beta(t)$. When the material contains time boundaries, the boundary conditions state that $\check{U}({\bf k},t)$ and $\check{\bf V}({\bf k},t)$ are continuous over those time boundaries. The second order wave equation (11) transforms to
@@ 6.1 Conservation of net field-momentum density
$\check{M}({\bf k},t)=\frac{1}{2}\Re(\check{U}^{*}\check{\bf V})\cdot{\bf n},$ (49)
@@
Propagation invariants in a time-dependent material are wave-field-related quantities that remain constant over time. We start by showing that for a source-free, time-dependent material, the net field-momentum density defined in equation (49) is a propagation invariant. To this end we analyze $\partial_{t}\check{M}({\bf k},t)=\frac{1}{4}\partial_{t}(\check{U}^{*}\check{\bf V}+\check{\bf V}^{*}\check{U})\cdot{\bf n}$. Applying the product rule for differentiation, using equations (41) and (42) for the source-free situation and the definition ${\bf n}={\bf k}/|{\bf k}|$, we obtain
$\partial_{t}\check{M}({\bf k},t)$ $=$ $\frac{1}{4}\Bigl{(}(\partial_{t}\check{U})^{*}\check{\bf V}+\check{U}^{*}\partial_{t}\check{\bf V}+(\partial_{t}\check{\bf V})^{*}\check{U}+\check{\bf V}^{*}\partial_{t}\check{U}\Bigr{)}\cdot{\bf k}/|{\bf k}|$
$=$ $\frac{1}{4}\biggl{(}\Bigl{(}\frac{1}{\beta}i{\bf k}\cdot\check{\bf V}^{*}\Bigr{)}\check{\bf V}-\check{U}^{*}\Bigl{(}\frac{1}{\alpha}i{\bf k}\check{U}\Bigr{)}+\Bigl{(}\frac{1}{\alpha}i{\bf k}\check{U}^{*}\Bigr{)}\check{U}-\check{\bf V}^{*}\Bigl{(}\frac{1}{\beta}i{\bf k}\cdot\check{\bf V}\Bigr{)}\biggr{)}\cdot{\bf k}/|{\bf k}|=0.$
@@
$\check{\jmath}({\bf k},t)=c^{2}(t)\check{M}({\bf k},t),$ (51)
which implies that for a time-dependent medium with piecewise continuous parameters the net power-flux density is not a propagation invariant, similar as noted in section 3.2 for $j({\bf x},t)$ over a single time boundary. Torrent et al. (2018) shows that amplitude amplification due to temporal variations can be used to compensate for energy loss due to material dissipation.
@@ 6.2 General propagation invariants
Next, we derive more general propagation invariants. We consider two mutually independent states (i.e., two solutions of equations (41) and (42)), which we distinguish with subscripts $A$ and $B$. We analyze the quantities $\partial_{t}(\check{U}_{A}\check{\bf V}_{B}-\check{\bf V}_{A}\check{U}_{B})\cdot(i{\bf k})$ and $\partial_{t}(\check{U}_{A}^{*}\check{\bf V}_{B}+\check{\bf V}_{A}^{*}\check{U}_{B})\cdot(i{\bf k})$. The latter can be seen as a generalization of $4\partial_{t}\check{M}({\bf k},t)=\partial_{t}(\check{U}^{*}\check{\bf V}+\check{\bf V}^{*}\check{U})\cdot{\bf n}$, with subscripts $A$ and $B$ added and ${\bf n}={\bf k}/|{\bf k}|$ replaced by $i{\bf k}$ (the latter replacement facilitates the inverse Fourier transformations in later sections). Applying the product rule for differentiation, using equations (41) and (42) for states $A$ and $B$, yields
@@ 7.2 Source-receiver reciprocity
This is the counterpart of the well-known source-receiver reciprocity relation ${\cal G}({\bf x}_{B},{\bf x}_{A},t,0)={\cal G}({\bf x}_{A},{\bf x}_{B},t,0)$ for an inhomogeneous time-independent medium, which states that the Green’s function between a source at ${\bf x}_{B}$ and a receiver at ${\bf x}_{A}$ is identical to the Green’s function between a source at ${\bf x}_{A}$ and a receiver at ${\bf x}_{B}$ (de Hoop, 1995; Morse & Feshbach, 1953). Equation (60) states that for a homogeneous time-dependent medium the causal Green’s function between a source at $t_{B}$ and a receiver at $t_{A}$ is identical to the acausal Green’s function between a sink at $t_{A}$ and a receiver at $t_{B}$. In the derivation we assumed that $t_{A}$ and $t_{B}$ lie both between $t_{b}$ and $t_{e}$. If we take $t_{b}\to-\infty$ and $t_{e}\to\infty$, then equations (58) – (60) hold for arbitrary $t_{A}$ and $t_{B}$. Note that for $t_{A}<t_{B}$ these expressions reduce to the trivial relation $0=0$.
@@ 7.3 Field-momentum density balance
$\check{M}({\bf k},t)\Bigr{|}_{t_{b}}^{t_{e}}$ $=$ $\frac{1}{2}\Re\int_{t_{b}}^{t_{e}}\bigl{(}\check{a}^{*}\check{\bf V}+\check{\bf b}^{*}\check{U}\bigr{)}\cdot{\bf n}{\rm d}t.$ (62)
@@ 8.2 Wave field representation for the case of a single time boundary
$\check{U}({\bf k},t^{\prime})=-\Bigl{[}\check{\cal G}_{2}({\bf k},{\bf 0},t^{\prime},t)\check{V}_{\rm d}({\bf k},t)+\beta_{2}\{\partial_{t}\check{\cal G}_{2}({\bf k},{\bf 0},t^{\prime},t)\}\check{U}({\bf k},t)\Bigr{]}_{t=t_{b}},\quad\mbox{for}\quad t^{\prime}>t_{b}.$ (69)
@@ 9 Green’s function retrieval
$\hskip 28.45274pt\beta(t_{b})\Bigl{[}\check{\cal G}^{*}({\bf k},{\bf 0},t_{A},t)\partial_{t}\check{\cal G}({\bf k},{\bf 0},t_{B},t)-\{\partial_{t}\check{\cal G}^{*}({\bf k},{\bf 0},t_{A},t)\}\check{\cal G}({\bf k},{\bf 0},t_{B},t)\Bigr{]}_{t=t_{b}}.$ (83)
@@ 10 Matrix-vector wave equation
The subscript ${\rm d}$ in ${\bf q}_{\rm d}$, ${\bf A}_{\rm d}$ and ${\bf d}_{\rm d}$ is used to distinguish these quantities from those in the 1D situation where, for example, wave field vector ${\bf q}$ contains the scalar wave fields $U$ and $V$ (Wapenaar et al., 2024). Equation (94) holds for a homogeneous time-dependent medium with continuous parameters $\alpha(t)$ and $\beta(t)$. When the material contains time boundaries, the boundary condition states that ${\bf q}_{\rm d}({\bf x},t)$ is continuous over those time boundaries. Equation (94) is the counterpart of a matrix-vector wave equation for an inhomogeneous time-independent material (Corones, 1975; Kosloff & Baysal, 1983; Fishman & McCoy, 1984; Ursin, 1983; Wapenaar & Berkhout, 1986; Løseth & Ursin, 2007).
@@ 11.1 Propagator matrix in space-time domain
As an illustration, we consider a piecewise constant material consisting of five time-independent slabs, with propagation velocities of 2000, 2000, 1200, 2500 and 1400 m/s, respectively. The parameter $\beta$ is taken constant. The duration of each time slab is 25 $\mu$s. We choose $t_{0}$ between the first and second time slab, i.e., $t_{0}=25\,\mu$s. Figure 4(a) shows a $z,t$-diagram of the propagator element $W_{\rm d}^{U,V}({\bf x},t,t_{0})$ (convolved with a spatial wavelet with a central wavenumber $k_{0}/2\pi=100$ m-1) for $x=0$ and $t_{0}=25\,\mu$s. The green line at $t_{0}=25\,\mu$s indicates the initial condition of equation (101), which for the considered off-diagonal element implies $W_{\rm d}^{U,V}({\bf x},t_{0},t_{0})=0$. In the first time slab right of this green line, the propagator consists (for $x=0$) of upward and downward propagating direct waves. At each time boundary (indicated by the dashed blue lines) these waves split into upward and downward propagating transmitted and reflected waves. This is a manifestation of the recursive character of the propagator, as formulated by equation (104).
Figure 4: Cross-sections at $x=0$ of (a) the propagator element $W_{\rm d}^{U,V}({\bf x},t,t_{0})$ and (b) the Green’s function ${\cal G}({\bf x},{\bf 0},t,t_{0})$ (both convolved with a spatial wavelet) for a piecewise constant time-dependent material (for Figure (a), see also movie Wuv.mp4 in the online material).
@@ 11.4 Relations between the Green’s function and the propagator matrix
Comparing this with equation (32) for ${\bf x}_{0}={\bf 0}$, with $c(t)$ defined in equation (12), we conclude that $H(t-t_{0})W_{\rm d}^{U,V}({\bf x},t,t_{0})$ indeed obeys the same wave equation as $-{\cal G}({\bf x},{\bf 0},t,t_{0})$. This completes the proof of equation (121). Figure 4(b) shows a $z,t$-diagram of the Green’s function ${\cal G}({\bf x},{\bf 0},t,t_{0})$ (convolved with a spatial wavelet) for $x=0$ and $t_{0}=25\,\mu$s. It has been obtained from $W_{\rm d}^{U,V}({\bf x},t,t_{0})$ in Figure 4(a) via equation (121). The green line at $t_{0}=25\,\mu$s indicates the causality condition of equation (33).
@@
Note that the simple relations between the Green’s function and the propagator matrix derived here are specific for a homogeneous time-dependent medium. They cannot be translated to the situation of an inhomogeneous time-independent medium by an exchange of space- and time-coordinates because the causality conditions for the Green’s function are the same for both types of medium (Wapenaar et al., 2024).
@@ 12 Conclusions
We used the reciprocity theorem of the space-convolution type to derive a source-receiver reciprocity relation for the Green’s function of a time-dependent material: the causal Green’s function between a source at $t_{B}$ and a receiver at $t_{A}$ is identical to the acausal Green’s function between a sink at $t_{A}$ and a receiver at $t_{B}$. We also used the reciprocity theorem of the space-convolution type to derive a general wave-field representation for a homogeneous time-dependent material. Similar as its counterpart for an inhomogeneous time-independent material, it forms the basis for the analysis of wave scattering problems. As an example, we used this representation to quantitatively explain the behaviour of the Green’s function after a single time boundary. We used the reciprocity theorem of the space-correlation type to derive a representation for Green’s function retrieval from passive measurements in a homogeneous, piecewise continuous time-dependent material.
@@
[…]
The analysis of wave propagation and scattering in materials that are simultaneously space- and time-dependent is not straightforward. Whereas effective medium theory can be used for periodic small-scale space-time variations (Trainiti & Ruzzene, 2016; Nassar et al., 2017b, a; Goldsberry et al., 2019; Sotoodehfar et al., 2023), the treatment of general piecewise continuous materials with mixed space-time boundaries and arbitrary varying space-time materials between these boundaries is more complex. One of the reasons is that $P$ and ${\bf Q}$ are the preferred wave field quantities to be analyzed at space boundaries and in space-dependent regions (case ii in section 2), whereas $U$ and ${\bf V}$ are the preferred quantities at time boundaries and in time-dependent regions (case iii in section 2). Many special situations are discussed in the literature. For example, Caloz & Deck-Léger (2020b) discusses how to model scattering at a mixed space-time boundary, Apffel & Fort (2022) shows how multiple distinct space and time boundaries can be used to control frequency conversion and Van Manen et al. (2024) discusses the design of an acoustic space-time material (with distinct space and time boundaries) which computes its own inverse. A fundamental treatment of wave propagation and scattering in arbitrary space-time materials remains subject to further research.
Data access
Links to movies associated with Figures 1, 3 and 4 and a matlab-code to reproduce the figures and movies can be found on https://gitlab.com/geophysicsdelft/OpenSource in the directory .../timematerial.
@@ Acknowledgements
The author thanks Johannes Aichele and Dirk-Jan van Manen for the inspiring discussions about waves in time-dependent materials. The constructive comments of the three referees are much appreciated and helped to improve the clarity of the paper.
@@ Appendix A Inverse Fourier Transform of wave-vector time domain Green’s function
The integral over $\theta$ is independent of $\phi$, so we may choose $\phi=0$. Furthermore, this integral from 00 to $2\pi$ is twice the real part of the integral from 00 to $\pi$. Hence
@@ Appendix B Propagator matrix for a time-independent slab
For the special case of a time-independent slab between $t_{n-1}$ and $t_{n}$, with parameters $\alpha_{n}$ and $\beta_{n}$, the solution of equation (105), with initial condition (106) (with $t_{0}$ replaced by $t_{n-1}$ and $t$ by $t_{n}$), reads
@@
with $\check{\cal G}({\bf k},{\bf 0},t_{n},t_{n-1})$ given in equation (47) for ${\bf x}_{0}={\bf 0}$, $t_{0}=t_{n-1}$ and $t=t_{n}$. Similarly, we obtain an explicit expression for $W_{\rm d}^{U,V}({\bf x},t_{n},t_{n-1})$ in the space-time domain by substituting equation (131) for ${\bf x}_{0}={\bf 0}$ into equation (126), using equation (60). This gives